A heart engine takes in 320 j of energy from a hot reservoir and performs 25.0 J of work in each cycle.find

(I). The efficiency of the engine and

(ii). The energy expelled to the cold reservoir in each cycle.


1
Expert's answer
2022-02-10T08:36:09-0500

Given:

"Q_1=\rm 320\: J"

"W=\rm 25.0\: J"


(i) The efficiency of the engine

"\eta=\frac{W}{Q_1}=\frac{25.0}{320}=0.078=7.8\%"

(ii) The energy expelled to the cold reservoir in each cycle

"Q_2=Q_1-W=320-25.0=295\:\rm J"